OpenStudy (lexiquesse):

f(x)=3x+2 g(x)=7x+6 find f*g and its domain

OpenStudy (lexiquesse):

is it 21x+12? and how do i find domain?

OpenStudy (sweetburger):

So your domain isnt restricted by any value of x.

OpenStudy (lexiquesse):

i'm not sure what you mean

OpenStudy (lexiquesse):

im guessing no

OpenStudy (sweetburger):

Do you remember FOIL?

OpenStudy (lexiquesse):

first outside inside last

OpenStudy (sweetburger):

Apply that to \[(3x+2)(7x+6)\]

OpenStudy (photon336):

domain means all possible x values for which the function is defined.

OpenStudy (lexiquesse):

21x^2+32x+12?

OpenStudy (sweetburger):

Yes that is the correct expansion.

OpenStudy (sweetburger):

Now both of the parent functions that you initially had were continuous on \[(-\infty,\infty)\] If you multiply two functions which are continuous generally you will result with a continuous function with a domain also being \[(-\infty,\infty)\]

OpenStudy (lexiquesse):

so the domain is -infinity to infinity?